The difference of two numbers is 18. Twice the smaller number is 56 less than three time the larger. What are the numbers?

Answer by checkley71(8403)

X-Y=18 OR X=18+Y
2Y=3X-56
2Y=3(18+Y)-56
2Y=54+3Y-56
2Y-3Y=-2
-Y=-2
Y=2 ANSWER
X-2=18
X=18+2
X=20 ANSWER.
PROOF
2*2=3*20-56
4=60-56
4=4